# Mitigating Visual Forgetting via Take-along Visual Conditioning for Multi-modal Long CoT Reasoning

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have demonstrated enhanced reasoning capabilities, evolving from Chain-of-Thought (CoT) prompting to advanced, product-oriented solutions like OpenAI o1. During our re-implementation of this model, we noticed that in multimodal tasks requiring visual input (e.g., geometry problems), Multimodal LLMs (MLLMs) struggle to maintain focus on the visual information, in other words, MLLMs suffer from a gradual decline in attention to visual information as reasoning progresses, causing text-over-relied outputs. To investigate this, we ablate image inputs during long-chain reasoning. Concretely, we truncate the reasoning process midway, then re-complete the reasoning process with the input image removed. We observe only a $\sim$2accuracy drop on MathVista’s test-hard subset, revealing the model’s textual outputs dominate the following reasoning process. Motivated by this, we propose Take-along Visual Conditioning (TVC), a strategy that shifts image input to critical reasoning stages and compresses redundant visual tokens via dynamic pruning. This methodology helps the model retain attention to the visual components throughout the reasoning. Our approach achieves state-of-the-art performance on average across five mathematical reasoning benchmarks (+3.4% vs previous sota), demonstrating the effectiveness of TVC in enhancing multimodal reasoning systems. The project page is available at https://sun-hailong.github.io/projects/TVC.
Machine Learning, ICML
@@ 2.1 Capturing the Visual Forgetting
As shown in Figure 2(a), we observe that at the 1/8 position of the reasoning process, the model effectively focuses on the visual inputs. However, as reasoning progresses, despite fluctuations in attention to visual tokens across layers, the model’s overall attention to visual evidence gradually decreases, leading to visual forgetting. Next, following the methodology of FastV (Chen et al., 2024a), we analyze the attention maps for several cases and find that the model predominantly focuses on previously generated text tokens rather than the input image. After approximately 20% tokens, the existence of image inputs on attention maps diminishes significantly, as illustrated in Figure 2(b). Both of the observations indicate a phenomenon of visual memory degradation, revealing the model’s limitations in maintaining consistent attention to visual inputs throughout the reasoning process.
@@ 3.1 Long-Chain Reasoning Data Collection
Prior research (Qin et al., 2024; Jiang et al., 2024) has identified two dominant paradigms for constructing long-chain reasoning systems: (1) explicit search-based methods, which utilize structures such as Monte Carlo Tree Search (MCTS) combined with specially trained reward models to guide the search process toward optimal solutions, and (2) instruction distillation approaches, which fine-tune models on curated long chain-of-thought (CoT) datasets. To efficiently develop an MLLM with long-chain reasoning capabilities, we adopt the distillation paradigm. In this section, we describe the distillation process and present the data generation pipeline aimed at enhancing the reasoning capability of MLLM.
Our distillation pipeline begins with aggregating publicly available reasoning datasets (e.g., MathV360K (Shi et al., 2024) and Geo170K (Gao et al., 2023)). Through an iterative process of distillation and response filtering, we progressively enhance the model’s long-chain reasoning capabilities. Specifically, the teacher model, QVQ-72B-Preview, generates long COT reasoning data, while the student model, Qwen2-VL, undergoes supervised fine-tuning on the filtered CoT data to enhance its reasoning performance.
@@ 4 Experiments
MiniCPM-V-2.6 (Yadav et al., 2025) 8B 60.8 18.4 17.6 9.8 - -
VITA-1.5 (Fu et al., 2025) 8B 66.2 19.5 23.4 9.6 - -
LLaVA-COT (Xu et al., 2024) 11B 52.5 19.9 22.6 7.8 - -
Qwen2-VL (Wang et al., 2024b) 7B 60.9 16.3 24.6 11.0 3.2 23.2
InternVL2.5 (Chen et al., 2024b) 8B 64.5 17.0 22.8 9.4 0.1 22.8
POINTS1.5 (Liu et al., 2024b) 8B 66.4 22.0 26.6 14.2 - -
Ovis1.6-Gemma2 (Lu et al., 2024b) 27B 70.2 20.6 37.8 17.0 - -
InternVL2.5-COT (Chen et al., 2024b) 78B 71.4 32.5 40.1 28.5 - -
LLaVA-OneVision (Li et al., 2024) 72B 67.1 25.3 27.2 15.6 - -
Qwen2-VL (Wang et al., 2024b) 72B 69.7 26.6 36.2 20.0 10.3 32.6
QVQ-72B-preview (QwenTeam, 2024) 72B 71.4 35.9 41.5 30.7 20.4 40.0
@@
We conducted comprehensive experiments across multiple vision-language benchmarks to demonstrate the effectiveness of our method. Section 4.1 provides implementation details for TVC. In Section 4.3, we present key results on visual reasoning tasks, supplemented with findings on general image understanding. Section 4.4 details ablation studies to evaluate the influence of critical design decisions. Finally, Section 4.5 presents visualizations and case studies to illustrate the method’s operational characteristics and insights derived from it.
@@ 4.1 Training Recipe
We integrated the TVC system with MLLMs of varying scales to demonstrate the effectiveness and generalizability of our approach. Initial implementation with Qwen2-VL-7B-Instruct confirmed the method’s validity. To further evaluate scalability and establish robust baselines against state-of-the-art MLLMs, we expanded the approach to a 72B model. Prior to training TVC, we followed the long-chain reasoning pipeline described earlier. We employed the LLaMA-Factory (Zheng et al., 2024) framework, using a learning rate of 2e-5, a batch size of 256, and 5 training epochs. During optimization, only the LLM parameters and cross-modal connector were trained, while the visual encoder remained frozen. The training process required 10 hours on a 64×H20 GPU setup for the 7B model and approximately 4 days for the 72B model. Additional details are presented in Table 4.
4.2 Evaluation Setup
We conduct a comprehensive experimental analysis across various visual reasoning benchmarks that require advanced visual reasoning skills. To ensure a well-rounded evaluation, we select several widely recognized and representative benchmarks, including MathVista (Lu et al., 2024a), MathVerse (Zhang et al., 2024a), MathVision (Wang et al., 2024a), Dynamath (Zou et al., 2024), and OlympiadBench (He et al., 2024). MathVista consists of 6,141 examples that require fine-grained, deep visual understanding and compositional reasoning. MathVerse contains 2,612 multi-subject math problems from a variety of sources. MathVision includes 3,040 high-quality mathematical problems sourced from established mathematics competitions. OlympiadBench features 8,476 bilingual multimodal problems tailored to Olympic-level mathematics and physics competitions. These benchmarks evaluate problem-solving abilities in mathematics, and following standard practice, we use GPT-4o-mini as the evaluator. Following the VLMEvalKit guidelines, we exclude the text-only split from MathVerse and the theorem-proof sections from OlympiadBench. For a fair comparison, we conduct evaluations using the testmini sets of MathVerse and MathVista. Fast evaluation is made possible through the use of the VLMEvalKit (Duan et al., 2024) and vLLM (Kwon et al., 2023).
@@ 4.3 Main Results on Visual Reasoning
Revised Paragraph: The results in Table 1 demonstrate the effectiveness and generalizability of TVCacross multiple visual reasoning benchmarks. Notably, our model achieves a 16.7% improvement over QVQ-72B-Preview on MathVision and a 17.6% gain on MathVerse, highlighting enhanced reasoning capabilities. Unlike conventional datasets where textual descriptions may include implicit visual cues, MathVerse is an all-around visual math benchmark specifically designed for equitable, in-depth evaluation of MLLMs. The significant gains on MathVerse underscore the significance of TVC, given the benchmark’s unique design principles. Furthermore, the TVC-7B model, despite its smaller size compared to counterparts, achieves competitive performance, even outperforming leading MLLMs in multiple cases. This demonstrates the robustness of our methodology even with more compact model configurations. Beyond task-specific visual reasoning, we extend our evaluation to general reasoning benchmarks (e.g., MathVista), where TVCconsistently delivers strong performance, achieving a 3.6% improvement over the original Qwen2-VL-72B model. These results emphasize TVC’s ability to excel in tasks requiring both perception and reasoning. Collectively, the findings indicate that TVC not only advances performance in specialized visual reasoning tasks but also offers substantial benefits in broader application scenarios.
@@ Appendix A Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) (Li et al., 2023; Liu et al., 2024a; Sun et al., 2024; Wang et al., 2024b; Lu et al., 2024b; McKinzie et al., 2024) integrate vision encoders (Radford et al., 2021; Zhai et al., 2023) with LLMs (AI, 2024; Bai et al., 2023), endowing them with robust capabilities across a wide range of domains. These include general visual understanding(OpenAI, 2024a; Li et al., 2024), mathematical reasoning (Shi et al., 2024; Gao et al., 2023), and answering college-level questions (Chen et al., 2023), demonstrating their versatility in real-world tasks. The rapid advancements in open-source models have also spurred the development of proprietary models, such as GPT-4o (OpenAI, 2024a), Gemini (Team et al., 2023; Reid et al., 2024), Qwen2-VL-MAX (Wang et al., 2024b), and Claude3 (anthropic, 2024). These models have demonstrated remarkable performance in both evaluation benchmarks and practical applications, solidifying their position at the forefront of AI research and deployment.
Reasoning with MLLMs. Recent advancements in MLLMs have significantly enhanced performance in reasoning tasks across both text and multimodal scenarios (OpenAI, 2024b; DeepSeek-AI, 2024; QwenTeam, 2024). Current methods primarily rely on CoT (Wei et al., 2022) to train MLLMs for step-by-step reasoning. Data-driven approaches include Math-LLaVA (Shi et al., 2024), which introduced the MathV360K dataset, and MAmmoTH-VL (Guo et al., 2024), which curates a large-scale multimodal CoT dataset in a scalable manner. Another line of research explores vision-text alignment. MAVIS (Zhang et al., 2024b) fine-tunes a math-specific vision encoder with curated caption data, while Math-PUMA (Zhuang et al., 2024) leverages the Kullback-Leibler (KL) divergence of next-token prediction distributions for modality alignment. In a different paradigm, MLLMs act as coordinators, utilizing external tools such as LLMs, web search engines, and computer programs for complex reasoning. Chameleon (Lu et al., 2023) orchestrates tool-call sequences, and Visual Sketchpad (Hu et al., 2024) enables models to generate visual sketches to aid reasoning.

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) have demonstrated enhanced reasoning capabilities, evolving from Chain-of-Thought (CoT) prompting to advanced, product-oriented solutions like OpenAI o1. During our re-implementation of this model, we noticed that in multimodal tasks requiring visual input (e.g., geometry problems), Multimodal LLMs (MLLMs) struggle to maintain focus on the visual information, in other words, MLLMs suffer from a gradual decline in attention to visual information as reasoning progresses, causing text-over-relied outputs. To investigate this, we ablate image inputs during long-chain reasoning. Concretely, we truncate the reasoning process midway, then re-complete the reasoning process with the input image removed. We observe only a $\sim$2accuracy drop on MathVista’s test-hard subset, revealing the model’s textual outputs dominate the following reasoning process. Motivated by this, we propose Take-along Visual Conditioning (TVC), a strategy that shifts image input to critical reasoning stages and compresses redundant visual tokens via dynamic pruning. This methodology helps the model retain attention to the visual components throughout the reasoning. Our approach achieves state-of-the-art performance on average across five mathematical reasoning benchmarks (+3.4 points vs previous sota), demonstrating the effectiveness of TVC in enhancing multimodal reasoning systems. The project page is available at https://sun-hailong.github.io/projects/TVC.
Mitigating Visual Forgetting via Take-along Visual Conditioning
for Multi-modal Long CoT Reasoning
Hai-Long Sun1,2 thanks: Work done during their internship/employment at Tencent Hunyuan, Zhun Sun4,5 11footnotemark: 1, Houwen Peng3, Han-Jia Ye1,2 thanks: Corresponding author: yehj@lamda.nju.edu.cn 1School of Artificial Intelligence, Nanjing University 2National Key Laboratory for Novel Software Technology, Nanjing University 3Tencent 4Center for Language AI Research, Tohoku University 5RIKEN Center for Advanced Intelligence Project
@@ 2.1 Capturing the Visual Forgetting
As shown in Figure 2(a), we observe that at the 1/8 position of the reasoning process, the model effectively focuses on the visual inputs. However, as reasoning progresses, despite fluctuations in attention to visual tokens across layers, the model’s overall attention to visual evidence gradually decreases, leading to visual forgetting. Next, following the methodology of FastV Chen et al. (2024a), we analyze the attention maps for several cases and find that the model predominantly focuses on previously generated text tokens rather than the input image. After approximately 20% tokens, the existence of image inputs on attention maps diminishes significantly, as illustrated in Figure 2(b). Both of the observations indicate a phenomenon of visual memory degradation, revealing the model’s limitations in maintaining consistent attention to visual inputs throughout the reasoning process.
@@ 3.1 Long-Chain Reasoning Data Collection
Prior research Qin et al. (2024); Jiang et al. (2024) has identified two dominant paradigms for constructing long-chain reasoning systems: (1) explicit search-based methods, which utilize structures such as Monte Carlo Tree Search (MCTS) combined with specially trained reward models to guide the search process toward optimal solutions, and (2) instruction distillation approaches, which fine-tune models on curated long chain-of-thought (CoT) datasets. To efficiently develop an MLLM with long-chain reasoning capabilities, we adopt the distillation paradigm. In this section, we describe the distillation process and present the data generation pipeline aimed at enhancing the reasoning capability of MLLM.
@@
Our distillation pipeline begins with aggregating publicly available reasoning datasets (e.g., MathV360K Shi et al. (2024) and Geo170K Gao et al. (2023)). Through an iterative process of distillation and response filtering, we progressively enhance the model’s long-chain reasoning capabilities. Specifically, the teacher model, QVQ-72B-Preview, generates long COT reasoning data, while the student model, Qwen2-VL, undergoes supervised fine-tuning on the filtered CoT data to enhance its reasoning performance.
@@ 4 Experiments
MiniCPM-V-2.6 Yadav et al. (2025) 8B 60.8 18.4 17.6 9.8 - -
VITA-1.5 Fu et al. (2025a) 8B 66.2 19.5 23.4 9.6 - -
LLaVA-COT Xu et al. (2024) 11B 52.5 19.9 22.6 7.8 - -
Qwen2-VL Wang et al. (2024b) 7B 60.9 16.3 24.6 11.0 3.2 23.2
InternVL2.5 Chen et al. (2024b) 8B 64.5 17.0 22.8 9.4 0.1 22.8
POINTS1.5 Liu et al. (2024b) 8B 66.4 22.0 26.6 14.2 - -
Ovis1.6-Gemma2 Lu et al. (2024b) 27B 70.2 20.6 37.8 17.0 - -
InternVL2.5-COT Chen et al. (2024b) 78B 71.4 32.5 40.1 28.5 - -
LLaVA-OneVision Li et al. (2024) 72B 67.1 25.3 27.2 15.6 - -
Qwen2-VL Wang et al. (2024b) 72B 69.7 26.6 36.2 20.0 10.3 32.6
QVQ-72B-preview QwenTeam (2024) 72B 71.4 35.9 41.5 30.7 20.4 40.0
@@
We conduct comprehensive experiments across multiple vision-language benchmarks to demonstrate the effectiveness of our method. Section 4.1 provides implementation details for TVC. In Section 4.3, we present key results on visual reasoning tasks, supplemented with findings on general image understanding. Section 4.4 details ablation studies to evaluate the influence of critical design decisions. Finally, Section 4.5 presents visualizations and case studies to illustrate the method’s operational characteristics and insights derived from it.
@@ 4.1 Training Recipe
We integrate the TVC system with MLLMs of varying scales to demonstrate the effectiveness and generalizability of our approach. Initial implementation with Qwen2-VL-7B-Instruct confirmed the method’s validity. To further evaluate scalability and establish robust baselines against state-of-the-art MLLMs, we expanded the approach to a 72B model. Prior to training TVC, we follow the long-chain reasoning pipeline described earlier. We use the LLaMA-Factory Zheng et al. (2024) framework, with a learning rate of 2e-5, a batch size of 256, and 5 training epochs. During optimization, only the LLM parameters and cross-modal connector are trained, while the visual encoder remains frozen. The training process requires 10 hours on a 64×H20 GPU setup for the 7B model and approximately 4 days for the 72B model. Additional details are presented in Table 4.
@@
4.2 Evaluation Setup
We conduct a comprehensive experimental analysis across various visual reasoning benchmarks that require advanced visual reasoning skills. To ensure a well-rounded evaluation, we select several widely recognized and representative benchmarks, including MathVista Lu et al. (2024a), MathVerse Zhang et al. (2024a), MathVision Wang et al. (2024a), Dynamath Zou et al. (2024), and OlympiadBench He et al. (2024). MathVista consists of 6,141 examples that require fine-grained, deep visual understanding and compositional reasoning. MathVerse contains 2,612 multi-subject math problems from a variety of sources. MathVision includes 3,040 high-quality mathematical problems sourced from established mathematics competitions. OlympiadBench features 8,476 bilingual multimodal problems tailored to Olympic-level mathematics and physics competitions. These benchmarks evaluate problem-solving abilities in mathematics, and following standard practice, we use GPT-4o-mini as the evaluator. Following the VLMEvalKit guidelines, we exclude the text-only split from MathVerse and the theorem-proof sections from OlympiadBench. For a fair comparison, we conduct evaluations using the testmini sets of MathVerse and MathVista. Fast evaluation is made possible through the use of the VLMEvalKit Duan et al. (2024) and vLLM Kwon et al. (2023).
@@ 4.3 Main Results on Visual Reasoning
The results in Table 1 demonstrate the effectiveness and generalizability of TVC across multiple visual reasoning benchmarks. Notably, our model achieves a 16.7% improvement over QVQ-72B-Preview on MathVision and a 17.6% gain on MathVerse, highlighting enhanced reasoning capabilities. Unlike conventional datasets where textual descriptions may include implicit visual cues, MathVerse is an all-around visual math benchmark specifically designed for equitable, in-depth evaluation of MLLMs. The significant gains on MathVerse underscore the significance of TVC, given the benchmark’s unique design principles. Furthermore, the TVC-7B model, despite its smaller size compared to counterparts, achieves competitive performance, even outperforming leading MLLMs in multiple cases. This demonstrates the robustness of our methodology even with more compact model configurations. Beyond task-specific visual reasoning, we extend our evaluation to general reasoning benchmarks (e.g., MathVista), where TVC consistently delivers strong performance, achieving a 3.6% improvement over the original Qwen2-VL-72B model. These results emphasize TVC’s ability to excel in tasks requiring both perception and reasoning. Collectively, the findings indicate that TVC not only advances performance in specialized visual reasoning tasks but also offers substantial benefits in broader application scenarios.
@@ 5 Conclusion
Acknowledgments
This work is partially supported by National Key R&D Program of China (2024YFE0202800), NSFC (62376118), Key Program of Jiangsu Science Foundation (BK20243012), CCF-Tencent Rhino-Bird Open Research Fund (RAGR20240101), AMED (JP25wm0625405), and Collaborative Innovation Center of Novel Software Technology and Industrialization.
@@ Appendix A Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) (Li et al., 2023; Liu et al., 2024a; Sun et al., 2025a; Wang et al., 2024b; Lu et al., 2024b; McKinzie et al., 2024; Sun et al., 2025c, b; Fu et al., 2025b; Dong et al., 2024) integrate vision encoders Radford et al. (2021); Zhai et al. (2023) with LLMs AI (2024); Bai et al. (2023), endowing them with robust capabilities across a wide range of domains. These include general visual understanding(OpenAI, 2024a; Li et al., 2024), mathematical reasoning (Shi et al., 2024; Gao et al., 2023), and answering college-level questions (Chen et al., 2023), demonstrating their versatility in real-world tasks. The rapid advancements in open-source models have also spurred the development of proprietary models, such as GPT-4o (OpenAI, 2024a), Gemini (Team et al., 2023; Reid et al., 2024), Qwen2-VL-MAX (Wang et al., 2024b), and Claude3 (anthropic, 2024). These models have demonstrated remarkable performance in both evaluation benchmarks and practical applications, solidifying their position at the forefront of AI research and deployment.
Reasoning with MLLMs. Recent advancements in MLLMs have significantly enhanced performance in reasoning tasks across both text and multimodal scenarios (OpenAI, 2024b; DeepSeek-AI, 2024; QwenTeam, 2024; Peng et al., 2025). Current methods primarily rely on CoT (Wei et al., 2022) to train MLLMs for step-by-step reasoning. Data-driven approaches include Math-LLaVA (Shi et al., 2024), which introduced the MathV360K dataset, and MAmmoTH-VL (Guo et al., 2024), which curates a large-scale multimodal CoT dataset in a scalable manner. Another line of research explores vision-text alignment. MAVIS (Zhang et al., 2024b) fine-tunes a math-specific vision encoder with curated caption data, while Math-PUMA (Zhuang et al., 2024) leverages the Kullback-Leibler (KL) divergence of next-token prediction distributions for modality alignment. In a different paradigm, MLLMs act as coordinators, utilizing external tools such as LLMs, web search engines, and computer programs for complex reasoning. Chameleon (Lu et al., 2023) orchestrates tool-call sequences, and Visual Sketchpad (Hu et al., 2024) enables models to generate visual sketches to aid reasoning.
@@ B.2 Reflection Word Pruning
Appendix C Discussion
However, as illustrated in Figure 8, we acknowledge several limitations. First, for highly complex reasoning tasks requiring sophisticated analytical capabilities, simply increasing visual revisits proves insufficient. In contrast, it is crucial to enhance the model’s inherent reasoning capacity. Second, our method assumes the availability of delayed visual processing, making it potentially unsuitable for real-time applications requiring instantaneous visual feedback, such as robotic navigation or time-sensitive decision-making scenarios.